OpenStudy (anonymous):

Linear equations. Solve for x. I will attempt to write it below as a formula. I have got several wrong and can't see what I did, please help and show steps.

OpenStudy (anonymous):

\[(x+5)/2 - (2+x)/3 = (2x + 1)/10\] Okay I can't see how to turn into a fraction-looking thing, so I used /.

OpenStudy (anonymous):

So, the first thing I did was try to times it all by 30 to get rid of denominators. I then got this: 15(x + 5) - 10(2 + x) = 3(2x + 1)

OpenStudy (lilg132):

15x + 75 - 20 - 10x = 6x + 3 5x + 55 = 6x + 3 55 - 3 = x 52 = x

OpenStudy (anonymous):

Oh that's cool actually the problem must be in the steps you wrote anyway because you got the right answer... I'm just looking now to see what I did wrong.

OpenStudy (anonymous):

Ah I put + 10x... there must be something about times-ing by -10 inside the brackets, I see. Sleep well and thanks.

OpenStudy (anonymous):

Yep. Just went through and all the mistakes were based around that error, eg -10(2x + 3) and I did 30 instead of -30.